a tennis ball is released from rest. (g= 9.8 m/s^2)
Calculate the distance it falls in 3 seconds.

calculate the velocity in 3 seconds.
Answer #1

According to motion equation

V=u+at

= 0+9.81×t=3*9.8=29.4m/s

So 29.4 will be object velocity after 3sec.